OpenStudy (anonymous):

Which of the following expressions are equivalent to −3.5(2 − 3n) − 2.5n? A. −7 − 8n B. −7 + 8n C. −7 − 13n D. −7 − n(10.5 − 2.5) E. −7 + n(10.5 − 2.5)

OpenStudy (anonymous):

Hmm,what do u think? @vazxlegend

OpenStudy (anonymous):

c

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

d

OpenStudy (anonymous):

No, there are actually 2 solutions..

OpenStudy (anonymous):

\[=-3.5(2-3n)-2.5n\]\[=-7+10.5n-2.5n\]\[=?\] solve it @vazxlegend

OpenStudy (anonymous):

a

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

c

OpenStudy (anonymous):

I thank so